### v1

# On the excitability of two-level atoms by spectrally encoded single-photon wave packets

Abstract
Single-photon wave packets are utilized as quantum information carriers within quantum information science and quantum communication applications. Here, we analyze the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet and obtain the excitation probability of the atom based on the Heisenberg-Langevin equations. Our results show that the spectral encoding process causes the single-photon wave packet to spread in the time domain, and the intensity of the wave packet decreases. Therefore, an encoded single-photon wave packet excites a two-level atom much weaker than an uncoded one. These results are crucial for understanding how encoded quantum light interacts with quantum nodes in realistic quantum networks. In particular, they reveal limitations and trade-offs in using spectrally encoded photons for secure, multiplexed quantum communication. The findings offer insights into optimizing encoding strategies for efficient atom-photon interactions and contribute to the development of scalable and secure quantum networking protocols.
keywords: Atom-photon interaction, Heisenberg-Langevin equations, Spectral encoding, Photon wave packet.
I Introduction
In recent years, much attention has been paid to the interaction between atom and light, which has many applications in quantum information, quantum communication, and fundamental experiments in quantum mechanics Lewenstein ; Rui ; Andersson ; Han ; Diaz ; Shomroni ; Tirandaz ; Moeini . In this regard, a single-photon wave packet is one of the most important candidates for examining atom-photon interaction in the experiments of quantum information processing Furusawa ; Brecht . Although quantum information can be encoded in different physical systems, such as trapped ions and superconducting devices, photonic states are much more practical because photons are indeed more robust against decoherence than many other quantum systems Flamini . Their ability to travel through optical fibers and free space with minimal loss makes them practical for carrying information between various nodes of a quantum network Gheri . Therefore, single-photon wave packets are good candidates for researchers in many fields, especially in quantum communication Hu , quantum computation Sheng2 , quantum key distribution Sych1 ; Sych2 , and entanglement Sheng1 .
Recently, many studies have examined the interaction between photon wave packets and atoms Slodicka ; Aljunid , molecules Wrigge , surface plasmons Chang and quantum dots Vamivakas . Furthermore, the interaction between a two-level atom and a single-photon wave packet has been extensively analyzed Lindlein ; Rephaeli ; Chen ; Rag ; Wang1 . However, the interaction of a two-level atom and an encoded single-photon wave packet has not been studied much.
In this paper, we analyze the time-dependent interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom which is in the ground state, based on the Heisenberg-Langevin equations. Then, we compare our results with the typical dynamic of interaction between a two-level atom and a single-photon wave packet reported in the literature Domokos ; Wang2 ; Tannoudji . We use the binary spectral encoding approach to generate the encoded single-photon wave packet provided in code division multiple-access (CDMA) communication systems Salehi .
The paper is organized as follows. In section II, we review a general formalism to analyze the dynamics of a two-level atom that interacts with a continuous mode state of quantum light. In section III, first, we introduce the spectral encoding approach. Then, we investigate the dynamics of interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom. Finally, the results are discussed in section IV.
II Model
A general formalism to study the dynamics of a two-level atom that interacts with a continuous mode state of quantum light is presented in Domokos ; Wang2 .
Let us consider the interaction of a two-level atom with the quantized photon wave packet as depicted in FIG. 1. The continuous mode state of a quantized field operator in the interaction picture is Loudon
$E(t)=i\int\operatorname{d}\omega(\frac{\hbar\omega}{4\pi\epsilon_{0}cA})^{1/2}\big{[}a(\omega)e^{-i\omega t}-a^{\dagger}(\omega)e^{i\omega t}\big{]}$ (1)
where $a^{\dagger}(\omega)$ and $a(\omega)$ are the creation and annihilation operators of the quantized field with the frequency $\omega$, respectively. The commutation relation of these operators is $\big{[}a(\omega),a^{\dagger}(\omega)\big{]}=\delta(\omega-\omega^{\prime})$, where $\delta$ denotes the Dirac delta function. Moreover, $\epsilon_{0}$, $c$ and $A$ are the vacuum permittivity, the speed of light and the cross section, respectively.
Figure 1: A schematic of the interaction between a two-level atom and a photon wave packet.
In the interaction picture and using rotating wave approximation (RWA), one can show that the interaction Hamiltonian of the system can be written as
$H_{\operatorname{I}}=-i\hbar\int\operatorname{d}\omega\lambda_{\omega}\big{[}\sigma_{+}a(\omega)e^{-i(\omega-\omega_{A})t}-\operatorname{H.C}\big{]}$ (2)
where $\omega_{A}$ is the atomic transition frequency and $\sigma_{z}=|e\rangle\langle e|-|g\rangle\langle g|$, $\sigma_{+}=|e\rangle\langle g|$, $\sigma_{-}=|g\rangle\langle e|$ are the atomic operators and H.C denotes Hermitian conjugates. Furthermore, the coupling strength is $\lambda_{\omega}=d\big{(}\omega/4\pi\epsilon_{0}\hbar cA\big{)}^{1/2}$ where $d$ is the value of the dipole moment of the atom.
One can show that the evolution of the atomic operators can be described by a set of Heisenberg-Langevin equations as Domokos ; Wang2 ; Tannoudji
$\frac{\operatorname{d}\sigma_{-}}{\operatorname{d}t}=-\gamma\sigma_{-}/2+\sigma_{z}\int\operatorname{d}\omega\lambda_{\omega}a(\omega)e^{-i(\omega-\omega_{A})t}+\zeta_{-}$ (3)
$\frac{\operatorname{d}\sigma_{z}}{\operatorname{d}t}=-\gamma(\sigma_{z}+1)-2\int\operatorname{d}\omega\lambda_{\omega}\big{[}\sigma_{+}a(\omega)e^{-i(\omega-\omega_{A})t}+\operatorname{H.C}\big{]}+\zeta_{z}$ (4)
where $\gamma$ is the decay rate of the atomic dipole and can be defined as $\gamma=2\pi\int\operatorname{d}\omega|\lambda_{\omega}|^{2}\delta(\omega-\omega_{A})$. Since we also considered the interaction between the atom and the environment, the terms of noise operators ($\zeta_{z}$ and $\zeta_{-}$) have appeared in the Heisenberg-Langevin equations, which are defined as
$\zeta_{-}=\sigma_{z}\int\operatorname{d}\omega^{\prime}\lambda_{\omega^{\prime}}b(\omega^{\prime})e^{-i(\omega^{\prime}-\omega_{A})t}$ (5)
$\zeta_{z}=$ $-2\int\operatorname{d}\omega^{\prime}\lambda_{\omega^{\prime}}\big{[}\sigma_{+}b(\omega^{\prime})e^{-i(\omega^{\prime}-\omega_{A})t}+\operatorname{H.C}\big{]}$ (6)
where $b^{\dagger}(\omega^{\prime})$ and $b(\omega^{\prime})$ are the creation and the annihilation operators of the environmental modes with the frequency $\omega^{\prime}$, respectively.
From now on, for simplicity, we use the narrowband approximation means that the spectral width of photon-wavepacket $\xi(\omega)$ is small compared to its central frequency $\omega_{0}$ Loudon . So, the coupling strength can be written as $\lambda_{\omega}=d\big{(}\omega_{0}/4\pi\epsilon_{0}\hbar cA\big{)}^{1/2}$.
II.1 Excitation of a two-level atom by a rectangular single-photon wave packet
We now calculate the excitation probability $P_{e}(t)$ of a two-level atom in interaction with a single-photon wave packet which can be defined as Scully
$P_{e}(t)=\frac{1}{2}\big{[}\langle\Psi_{0}|\sigma_{z}(t)|\Psi_{0}\rangle+1\big{]}$ (7)
We suppose the initial state of the total system is $|\Psi_{0}\rangle=|g\rangle|1_{\xi}\rangle|0_{e}\rangle$ where is a product state of the atomic ground state, the single-photon wave packet state, and the vacuum state of the environment, respectively.
Figure 2: The excitation probability $P_{e}(t)$ of a two-level atom that interacts with a rectangular shape of a single-photon wave packet as a function of $\gamma t$.
A single-photon wavepacket state can be defined as $|1_{\xi}\rangle=a^{\dagger}_{\xi}|0\rangle$ where
$a^{\dagger}_{\xi}=\int\operatorname{d}\omega\xi(\omega)a^{\dagger}(\omega)$ (8)
where $\xi(\omega)$ and $a^{\dagger}_{\xi}$ represent the spectral amplitude of the photon wave packet and the photon wave packet creation operator, respectively Loudon . Moreover, we assume the environment to be initially in the vacuum state. Based on the Born approximation, the changes in the state of the environment are negligible Schlosshauer . So, the state of the environment remains vacuum at all times. Considering the initial vacuum state of the environment, the average values of the noise operators in Eq. (3) and Eq.(4) will be zero.
To obtain the excitation probability $P_{e}(t)$ of a two-level atom, we need to calculate the expectation value of the atomic operator $\sigma_{z}$ as $\langle\sigma_{z}(t)\rangle$ in Eq. (7). So, one should solve the set of coupled differential equations for the atomic operators ($\sigma_{z}$ and $\sigma_{\pm}$) in Eq. (3) and Eq. (4). Moreover, we should define the explicit shape of the single-photon wave packet.
In this regard, we take a single-photon wave packet with a rectangular shape as an example Salehi :
$\xi(\omega)=\begin{cases}\sqrt{P_{0}}/W&\mbox{for}-\frac{W}{2}\leq\omega\leq\frac{W}{2}\\ 0&\mbox{elsewhere}\end{cases}$ (9)
where $P_{0}$ and $W$ are the peak power and the total bandwidth of the pulse, respectively. Therefore, one can show that the temporal shape pulse is a sinc function:
$\xi(t)=\sqrt{P_{0}}\mbox{sinc}(\frac{W}{2}t)$ (10)
where sinc x=sin $\mbox{sinc x}=\frac{\mbox{sin x}}{\mbox{x}}$.
Assuming $W=1.5\gamma$, we can obtain the excitation probability $P_{e}(t)$ of a two-level atom that interacts with the rectangular shape of a single-photon wave packet. In FIG. 2, we have plotted the excitation probability $P_{e}(t)$ of the atom as a function of $\gamma t$. This figure shows the damping behavior of the atom in the presence of the environment.
III Encoded single-photon wave packets
In this section, we study the interaction between a two-level atom and an encoded single-photon wave packet. First, we review the spectral phase encoding approach.
III.1 Spectral encoding approach
The spectral phase encoding approach to encode the spectrum of a photon wave packet is used in CDMA communication systems. See Salehi for a general review.
A spectral phase-shifting operator $U$ with a phase value $\theta$ can be written as
$U=e^{-i\sum_{\omega}\theta(\omega)a^{\dagger}(\omega)a(\omega)}=\prod_{\omega}U(\omega)$ (11)
where $\theta(\omega)$ is the phase shift of the frequency $\omega$ on the spectral wave packet of the pulse $\xi(\omega)$ and $U(\omega)=e^{-i\theta(\omega)a^{\dagger}(\omega)a(\omega)}$. In the Heisenberg picture, one can show that:
$Ua^{\dagger}(\omega)U^{\dagger}=a^{\dagger}(\omega)e^{-i\theta(\omega)}$ (12)
Applying Eq. (12) to the photon wave packet creation operator in Eq. (8), we have $Ua^{\dagger}_{\xi}U^{\dagger}=a^{\dagger}_{\xi^{e}}$ where
$\xi^{e}(\omega)=\xi(\omega)e^{-i\theta(\omega)}$ (13)
in which superscript $e$ denotes the encoded photon wave packet. Moreover, the encoded wave packet in the temporal domain can be obtained by the Fourier transform as
$\xi^{e}(t)=\frac{1}{\sqrt{2\pi}}\int\operatorname{d}\omega\xi(\omega)e^{-i\big{(}\omega t+\theta(\omega)\big{)}}$ (14)
Therefore, we conclude that the spectral encoding process just affects the spectral amplitude $\xi(\omega)$ of the quantum wave packet Salehi .
III.2 Excitation of a two-level atom by an encoded single-photon wave packet
To encode the single-photon wave packet with a rectangular shape in Eq. (10), we divided the spectrum of the wave packet $\xi(\omega)$ into $N_{0}$ sequential and distinct frequency chips, each bandwidth is $\Omega=W/N_{0}$. Based on the spectral encoding approach, one can show that the time domain representation of the encoded rectangular wave packet can be written as Salehi
$\xi^{e}(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}\sum_{n=-N}^{N}\mbox{exp}\big{[}-i(n\Omega t)+\phi_{n}\big{]}$ (15)
where the phase $\phi_{n}$ denotes the $n$th element of a code consisting of $N_{0}=2N+1$. Moreover, the term $\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}$ is a real function that represents the temporal width of the encoded wave packet and the exponential term shows a periodic behavior with period $T=2\pi/\Omega$ which its temporal shape depends on the code elements $\phi_{n}$. Note that if we have $\phi_{n}=0$ for all $n$, Eq. (15) reduces to Eq. (10) Salehi .
We now consider $N_{0}=3,5,7,31,63$ for the code length of an encoded single-photon wave packet to calculate the excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded pulse according to Eq. (7). In addition, we assume the value of the phase $\phi_{n}$ is 00 and $\pi$ which correspond to the amplitude of $+1$ and $-1$, respectively. This spectral encoding process is called random binary spectral encoding Salehi .
We have plotted the excitation probability $P_{e}(t)$ of a two-level atom for the various code lengths of the encoded single-photon wave packet, which is compared to the uncoded wave packet in FIG. 3. We conclude that the encoding process of the single-photon wave packet decreases the excitation probability of a two-level atom. To analyze this result, we have plotted the intensity of the uncoded and encoded wave packet in FIG. 4. As is clear in this figure, the encoding process causes the single-photon wave packet to spread in time domain. Consequently, the power of the wave packet diminishes, and the wave packet behaves as a low-intensity pulse. Therefore, an encoded wave packet excites a two-level atom much weaker than an uncoded one.
Figure 3: The excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded rectangular shape of photon wave packet as a function of $\gamma t$ for the various code length $N_{0}$ (blue curves) compared to the uncoded case (red curves). a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
Figure 4: Intensity of the uncoded (orange) and encoded (blue) single-photon wave packet. In this figure, $N_{0}=31$.
The question that arises here is: what is the origin of the oscillations in the excitation probability of the atom that interacts with an encoded photon wave packet?. To answer this question, we investigate the relationship between the excitation probability of the atom and the sequences of the random phase (0 and $\pi$) in the spectral encoding process according to Eq. (15). In FIG. 5, we have plotted the excitation probability of the atom considering three different sequences of the random phase for the code length $N_{0}=63$. Our results show that the oscillations in the excitation probability of the atom depend on the selected sequences of the random phase in the binary spectral encoding process.
Figure 5: The excitation probability $P_{e}(t)$ of a two-level atom by an encoded photon wave packet considering three different sequences of the random phase (0 and $\pi$) for the code length $N_{0}=63$ in the binary spectral encoding process.
We also analyze the effect of the code length of the encoded photon wave packet on the excitation probability of a two-level atom. We have plotted the relationship between the excitation probability $P_{e}(t)$ of the atom and the code length $N_{0}$ of the encoded single-photon wave packet in FIG. 6. This figure shows that the excitation probability decreases with increasing code length. Since the code length increases, the photon wave packet spreading rate increases significantly (see FIG. 7). Consequently, the ability of the encoded wave packet to excite the atom decreases.
Figure 6: The excitation probability $P_{e}(t)$ of a two-level atom as a function of the code length $N_{0}$ of the encoded single-photon wave packet.
Figure 7: Spreading the encoded photon wave packet in time domain for the various code lengths. a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
IV Conclusion
We studied the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet based on the Heisenberg-Langevin equations. We used a spectral phase encoding method inspired by classical CDMA systems to encode quantum information in the frequency domain of the photon. Considering this formalism, we showed that for an encoded wave packet, the excitation probability of the atom is much lower than an unencoded one. On the other hand, the encoding process causes the wave packet to spread in the time domain and the ability to excite a two-level atom decreases. Moreover, we evaluated how encoding parameters—such as code length and phase randomness—affect atomic excitation probabilities.
These findings are particularly relevant for quantum networks, where atoms serve as quantum nodes and qubits are transmitted over long distances via single-photon wave packets. In such systems, an encoded single-photon wave packet carries quantum information to the receiver. If the receiver applies a decoding mask that is the conjugate of the original encoding mask, the photon’s spectral profile is correctly reconstructed, allowing it to interact effectively with the atomic node. However, if the decoding mask does not match the encoding mask, the wave packet fails to recover its original shape and remains a temporally spread, low-intensity pulse. As a result, it has minimal impact on the atomic excitation. This behavior provides a built-in filtering mechanism: atoms can act as selective photodetectors, responding only to correctly encoded photons while rejecting mismatched or unauthorized signals. Such properties are valuable for enhancing the security and robustness of quantum communication protocols.
Moreover, this research provides a fundamental insight into how spectral encoding, commonly used in CDMA systems, affects the quantum mechanical behavior of light-matter interaction. Understanding this effect is essential for developing robust quantum communication protocols that rely on encoding information in the frequency domain. Our findings indicate that while spectral encoding enhances information security and multiplexing capabilities, it also introduces trade-offs in terms of interaction efficiency with quantum nodes, which must be addressed in the system design. This study also opens the door to future research in optimizing spectral encoding schemes that balance encoding complexity and atom-photon interaction strength. In practical implementations of quantum repeaters or quantum routers, such trade-offs must be well-understood to ensure high fidelity and reliability of quantum information transfer. Furthermore, the methodology presented here can be extended to investigate other encoding formats, multi-level atoms, and more complex photonic structures, including entangled wave packets or time-bin encodings.
Ultimately, our work contributes to the theoretical foundation necessary for advancing quantum network architectures and can inform experimental setups aiming to test photon-mediated interactions at the quantum level with encoded information. The ability to predict and control atomic excitation probabilities in the presence of encoded quantum signals is a step toward realizing scalable and secure quantum communication infrastructures.
Acknowledgment
The author would like to thank Jawad Salehi, Mohammad Rezai, and Ali Soltanmanesh for their instructive comments and useful discussion.
Data Availability
The data that support the findings of this study are available on request from the corresponding author

### v2

@@
-On the excitability of two-level atoms by spectrally encoded single-photon wave packets
+On the excitability of two-level atoms by spectrally encoded single-photon wave packets in quantum networks
 Abstract
-Single-photon wave packets are utilized as quantum information carriers within quantum information science and quantum communication applications.
-Here, we analyze the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet and obtain the excitation probability of the atom based on the Heisenberg-Langevin equations.
-Our results show that the spectral encoding process causes the single-photon wave packet to spread in the time domain, and the intensity of the wave packet decreases.
-Therefore, an encoded single-photon wave packet excites a two-level atom much weaker than an uncoded one.
-These results are crucial for understanding how encoded quantum light interacts with quantum nodes in realistic quantum networks.
-In particular, they reveal limitations and trade-offs in using spectrally encoded photons for secure, multiplexed quantum communication.
-The findings offer insights into optimizing encoding strategies for efficient atom-photon interactions and contribute to the development of scalable and secure quantum networking protocols.
-keywords: Atom-photon interaction, Heisenberg-Langevin equations, Spectral encoding, Photon wave packet.
+We analyze the time-dependent interaction between a two-level atom and a spectrally encoded single-photon wave packet using the Heisenberg-Langevin equations and derive the atomic excitation probability.
+Spectral phase encoding broadens the photon wave packet in the time domain and reduces its peak intensity, leading to markedly weaker atomic excitation than for an unencoded photon.
+We formalize this behavior via an overlap bound with the time-reversed spontaneous-emission mode and show how excitation depends on code length, bandwidth, and phase errors.
+Interpreted at the quantum network level, atoms behave as phase-sensitive, and mode-selective receivers whose response scales with a spectral-overlap functional that captures decoding fidelity, detuning, and multiuser interference.
+From this, we extract design rules and performance bounds for encoded links, quantifying trade-offs among code length, addressability, cross-talk, and identifying tolerances for decoding error.
+These results clarify how spectrally encoded photons couple to quantum nodes and provide guidelines for efficient, scalable, and secure quantum networking.
+keywords: Atom-photon interaction, Heisenberg-Langevin equations, Spectral encoding, Quantum networks.
 I Introduction
@@
 We use the binary spectral encoding approach to generate the encoded single-photon wave packet provided in code division multiple-access (CDMA) communication systems Salehi .
+Moreover, as we know, a central challenge for quantum networks is the selective addressability of remote atomic or solid-state nodes while sharing photonic resources among many users.
+In this context, spectral encoding provides a high-dimensional multiplexing scheme compatible with optical fibers and ultrafast pulse shaping.
+However, atom-photon interfaces are inherently mode-selective, as the coupling efficiency depends on the spectral-temporal overlap between the photon wave packet and the atomic or cavity response.
+In this work, we quantify this overlap for encoded single-photon wave packets and, crucially, translate it into network-level figures of merit-decoding fidelity, multiuser cross-talk, and mismatch tolerance, thereby turning our microscopic model into concrete design guidance for scalable, multiplexed quantum links.
 The paper is organized as follows.
@@
 Then, we investigate the dynamics of interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom.
-Finally, the results are discussed in section IV.
+In section IV, we analyze the applications of our approach in quantum networks.
+Finally, the results are discussed in section V.
 II Model
@@
 In this regard, we take a single-photon wave packet with a rectangular shape as an example Salehi :
-$\xi(\omega)=\begin{cases}\sqrt{P_{0}}/W&\mbox{for}-\frac{W}{2}\leq\omega\leq\frac{W}{2}\\ 0&\mbox{elsewhere}\end{cases}$ (9)
+[…]
 where $P_{0}$ and $W$ are the peak power and the total bandwidth of the pulse, respectively. Therefore, one can show that the temporal shape pulse is a sinc function:
 $\xi(t)=\sqrt{P_{0}}\mbox{sinc}(\frac{W}{2}t)$ (10)
-where sinc x=sin $\mbox{sinc x}=\frac{\mbox{sin x}}{\mbox{x}}$.
+[…]
 Assuming $W=1.5\gamma$, we can obtain the excitation probability $P_{e}(t)$ of a two-level atom that interacts with the rectangular shape of a single-photon wave packet.
@@
 This figure shows the damping behavior of the atom in the presence of the environment.
+II.2 Optimal temporal mode and an overlap bound
+It is well known that, for a two-level atom coupled to a propagating mode under the Markov approximation, the time-reversed spontaneous-emission mode maximizes single-photon absorption Stobinska .
+Denoting the optimal (normalized) temporal profile by
+$\xi_{\operatorname{opt}}(t)=\sqrt{\gamma}e^{\gamma t/2}\Theta(-t)$ (11)
+where $\Theta(-t)$ is a Heaviside step function, the excitation probability achieved by an arbitrary normalized photon temporal mode $\xi(t)$ obeys the overlap bound
+$P_{e}^{\max}\leq\beta|\int_{-\infty}^{\infty}\operatorname{d}t\xi^{*}(t)\xi_{\operatorname{opt}}(t)|^{2}=\beta|\langle\xi\mid\xi_{\operatorname{opt}}\rangle|^{2}$ (12)
+where $\beta\in[0,1]$ denotes the coupling efficiency into the atomic radiation mode for free-space focusing or waveguide coupling.
+In other words, $\beta$ measures the fraction of the atom’s emission that goes into the mode of interest-equivalently, the mode-matching efficiency or branching ratio into that channel.
+Moreover, we have $\int\operatorname{d}t\mid\xi(t)\mid^{2}=1$.
+The bound in Eq.
+(12) follows directly from the Heisenberg-Langevin Eqs.
+(3)–(4) by writing the atomic drive as a convolution with the impulse response $h(t)=\sqrt{\gamma}e^{-\gamma t/2}\Theta(t)$ and using Cauchy–Schwarz, the equality is saturated for $\xi=\xi_{\operatorname{opt}}$.
+In our simulations, $\beta=1$ (ideal coupling), so Eq.
+(12) provides a tight, parameter-free reference for all cases studied.
+The derivation details of Eq.
+(12) are given in Appendix A.
 III Encoded single-photon wave packets
@@
 Based on the spectral encoding approach, one can show that the time domain representation of the encoded rectangular wave packet can be written as Salehi
-$\xi^{e}(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}\sum_{n=-N}^{N}\mbox{exp}\big{[}-i(n\Omega t)+\phi_{n}\big{]}$ (15)
+$\xi^{e}(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}\big(\frac{\Omega}{2}t\big)\sum_{n=-N}^{N}\mbox{exp}\big[-i(n\Omega t+\phi_{n})\big]$ (17)
 where the phase $\phi_{n}$ denotes the $n$th element of a code consisting of $N_{0}=2N+1$.
-Moreover, the term $\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}$ is a real function that represents the temporal width of the encoded wave packet and the exponential term shows a periodic behavior with period $T=2\pi/\Omega$ which its temporal shape depends on the code elements $\phi_{n}$.
+Moreover, the term $\mbox{sinc}\big(\frac{\Omega}{2}t\big)$ is a real function that represents the temporal width of the encoded wave packet and the exponential term shows a periodic behavior with period $T=2\pi/\Omega$ whose temporal shape depends on the code elements $\phi_{n}$.
 Note that if we have $\phi_{n}=0$ for all $n$, Eq.
 (17) reduces to Eq.
-(10) Salehi .
+(10).
+Figure 3: The excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded rectangular shape of photon wave packet as a function of $\gamma t$ for the various code length $N_{0}$ (blue curves) compared to the uncoded case (red curves).
+a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
 We now consider $N_{0}=3,5,7,31,63$ for the code length of an encoded single-photon wave packet to calculate the excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded pulse according to Eq.
 (7).
-In addition, we assume the value of the phase $\phi_{n}$ is 00 and $\pi$ which correspond to the amplitude of $+1$ and $-1$, respectively.
+In addition, we assume the value of the phase $\phi_{n}$ is 0 and $\pi$ which correspond to the amplitude of +1+1 and 1-1, respectively.
 This spectral encoding process is called random binary spectral encoding Salehi .
+In the binary spectral encoding, we intentionally choose odd $N_{0}$ so that a central chip is centered at $\omega_{0}$ (the atomic transition), yielding a symmetric index set $n\in\{-N,...,0,...,N\}$.
+This choice avoids placing $\omega_{0}$ in between two chips (the even-$N_{0}$ case), which tends to introduce a systematic $\cos(\Omega t/2)$ modulation in the time-domain envelope and can lead to strong destructive interference at $t=0$ for many random $\{0,\pi\}$.
+We verified numerically that even $N_{0}$ exhibits a larger fraction of near-vanishing overlaps with $\xi_{\operatorname{opt}}(t)$, consistent with the analytical discussion below.
+In other words, for even $N_{0}$, the resonance lies between the two central chips $n\pm\frac{1}{2}$.
+Writing the sum over half-integer indices introduces an extra factor $\cos(\Omega t/2)$ in $c_{\phi}(t)$, which suppresses the integrand near $t\approx 0$ for random $\{0,\pi\}$ phases and thus yields a larger set of near-zero overlaps in Eq.
+(12).
 We have plotted the excitation probability $P_{e}(t)$ of a two-level atom for the various code lengths of the encoded single-photon wave packet, which is compared to the uncoded wave packet in FIG.
@@
 Therefore, an encoded wave packet excites a two-level atom much weaker than an uncoded one.
-Figure 3: The excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded rectangular shape of photon wave packet as a function of $\gamma t$ for the various code length $N_{0}$ (blue curves) compared to the uncoded case (red curves).
-a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
-Figure 4: Intensity of the uncoded (orange) and encoded (blue) single-photon wave packet. In this figure, $N_{0}=31$.
 The question that arises here is: what is the origin of the oscillations in the excitation probability of the atom that interacts with an encoded photon wave packet?.
@@
 Our results show that the oscillations in the excitation probability of the atom depend on the selected sequences of the random phase in the binary spectral encoding process.
-Figure 5: The excitation probability $P_{e}(t)$ of a two-level atom by an encoded photon wave packet considering three different sequences of the random phase (0 and $\pi$) for the code length $N_{0}=63$ in the binary spectral encoding process.
+Figure 4: Intensity of the uncoded (orange) and encoded (blue) single-photon wave packet. In this figure, $N_{0}=31$.
+Figure 5: The excitation probability $P_{e}(t)$ of a two-level atom by an encoded photon wave packet considering three different sequences of the random phase (0 and $\pi$) for the code length $N_{0}=63$ in the binary spectral encoding.
 We also analyze the effect of the code length of the encoded photon wave packet on the excitation probability of a two-level atom.
@@
 Figure 6: The excitation probability $P_{e}(t)$ of a two-level atom as a function of the code length $N_{0}$ of the encoded single-photon wave packet.
-Figure 7: Spreading the encoded photon wave packet in time domain for the various code lengths. a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
-IV Conclusion
-We studied the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet based on the Heisenberg-Langevin equations.
-We used a spectral phase encoding method inspired by classical CDMA systems to encode quantum information in the frequency domain of the photon.
-Considering this formalism, we showed that for an encoded wave packet, the excitation probability of the atom is much lower than an unencoded one.
-On the other hand, the encoding process causes the wave packet to spread in the time domain and the ability to excite a two-level atom decreases.
-Moreover, we evaluated how encoding parameters—such as code length and phase randomness—affect atomic excitation probabilities.
-These findings are particularly relevant for quantum networks, where atoms serve as quantum nodes and qubits are transmitted over long distances via single-photon wave packets.
-In such systems, an encoded single-photon wave packet carries quantum information to the receiver.
-If the receiver applies a decoding mask that is the conjugate of the original encoding mask, the photon’s spectral profile is correctly reconstructed, allowing it to interact effectively with the atomic node.
-However, if the decoding mask does not match the encoding mask, the wave packet fails to recover its original shape and remains a temporally spread, low-intensity pulse.
-As a result, it has minimal impact on the atomic excitation.
-This behavior provides a built-in filtering mechanism: atoms can act as selective photodetectors, responding only to correctly encoded photons while rejecting mismatched or unauthorized signals.
-Such properties are valuable for enhancing the security and robustness of quantum communication protocols.
-Moreover, this research provides a fundamental insight into how spectral encoding, commonly used in CDMA systems, affects the quantum mechanical behavior of light-matter interaction.
-Understanding this effect is essential for developing robust quantum communication protocols that rely on encoding information in the frequency domain.
-Our findings indicate that while spectral encoding enhances information security and multiplexing capabilities, it also introduces trade-offs in terms of interaction efficiency with quantum nodes, which must be addressed in the system design.
-This study also opens the door to future research in optimizing spectral encoding schemes that balance encoding complexity and atom-photon interaction strength.
-In practical implementations of quantum repeaters or quantum routers, such trade-offs must be well-understood to ensure high fidelity and reliability of quantum information transfer.
-Furthermore, the methodology presented here can be extended to investigate other encoding formats, multi-level atoms, and more complex photonic structures, including entangled wave packets or time-bin encodings.
-Ultimately, our work contributes to the theoretical foundation necessary for advancing quantum network architectures and can inform experimental setups aiming to test photon-mediated interactions at the quantum level with encoded information.
-The ability to predict and control atomic excitation probabilities in the presence of encoded quantum signals is a step toward realizing scalable and secure quantum communication infrastructures.
+Figure 7: Spreading the encoded photon wave packet in the time domain for the various code lengths. a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
+Using Eq. (17), the normalized encoded temporal profile can be expressed as a product of a slowly varying envelope $s(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}(\frac{\Omega t}{2})$ and a coded comb
+$c_{\phi}(t)=\sum_{n=-N}^{N}\exp[-in\Omega t+i\phi_{n}]$ (18)
+so that $\xi_{e}(t)\propto s(t)c_{\phi}(t)$. The overlap in Eq. (12) becomes
+$\langle\xi_{e}\mid\xi_{\operatorname{opt}}\rangle\propto\int\operatorname{d}t\mbox{sinc}(\frac{\Omega t}{2})e^{\gamma t/2}\Theta(-t)\sum_{n=-N}^{N}\exp[in\Omega t-i\phi_{n}]$ (19)
+Evaluating the time integral term-wise yields a closed-form dependence on $\Omega/\gamma$ and on the code phases $\{\phi_{n}\}$ via a weighted discrete Fourier transform of the sequence $\{e^{-i\phi_{n}}\}$.
+Consequently, (i) increasing the code length $N_{0}$ broadens $c_{\phi}(t)$ and reduces the overlap, explaining the monotonic trend of FIG.
+6; and (ii) different phase sequences produce distinct beating patterns in $P_{e}$ harmonics of $\Omega$ consistent with the oscillations seen in FIG.
+5.
+IV Applications to quantum networks: mode-selective atom-photon interfaces with spectral encoding
+IV.1 Decoding as spectral-phase compensation
+Let the encoder impose a phase mask $\theta(\omega)$, producing the encoded spectrum $\xi^{e}(\omega)=\xi(\omega)e^{-i\theta(\omega)}$.
+A receiver implements a decoding mask $D(\omega)=e^{i\tilde{\theta}(\omega)}$ (e.g., via a $4f$ pulse shaper), yielding an effective input $\xi_{in}(\omega)=\xi(\omega)e^{i[\tilde{\theta}(\omega)-\theta(\omega)]}$.
+We define the residual phase error $\phi(\omega)=\tilde{\theta}(\omega)-\theta(\omega)$.
+So, perfect decoding corresponds to $\phi(\omega)=0$.
+The atomic response in the frequency domain is well captured by the Lorentzian susceptibility (narrowband approximation in Section II) Scully
+$\chi_{A}(\omega)=\frac{\sqrt{\gamma}}{\gamma/2-i(\omega-\omega_{A})}$ (20)
+where $\gamma$ is the dipole decay rate and $\omega_{A}$ is the atomic transition frequency.
+In the single-photon, weak-excitation regime used throughout, the excitation amplitude is proportional to the scalar product of the input mode with $\chi_{A}$
+$\mathcal{A}\propto\int\operatorname{d}\omega\xi(\omega)e^{i\phi(\omega)}\chi_{A}(\omega)$ (21)
+We define the spectral-overlap functional Fiechtner
+$\mathcal{M}[\phi,\Delta]=\frac{\int\operatorname{d}\omega\xi(\omega)e^{i\phi(\omega)}\chi_{A}(\omega-\Delta)}{\Big(\int\operatorname{d}\omega\mid\xi(\omega)\mid^{2}\Big)^{1/2}\Big(\int\operatorname{d}\omega\mid\chi_{A}(\omega)\mid^{2}\Big)^{1/2}}$ (22)
+with dtuning $\Delta=\omega-\omega_{A}$.
+The spectral-overlap functional was defined by inserting the photon spectrum $\xi(\omega)$ and the atomic susceptibility $\chi_{A}(\omega)$ into the excitation amplitude integral and normalizing it to be dimensionless.
+It is simply the normalized inner product between the photon mode and the atomic response.
+This functional directly measures how well the photon’s spectral-temporal mode matches the atomic absorption mode, and the excitation probability satisfies $P_{e}^{\max}\propto\mathcal{M}[\phi,\Delta]\mid^{2}$.
+Then, the peak excitation probability obeys the bound in Eq.
+(12)
+$P_{e}^{\max}\leq\beta\kappa\mid\mathcal{M}[\phi,\Delta]\mid^{2}$ (23)
+where $\beta\in[0,1]$ is the coupling efficiency into the atomic mode (free-space or waveguide) and $\kappa$ is a geometry-dependent constant (absorbed in our normalization in Section II).
+Equality is approached when temporal mode matching is optimal.
+This recovers our observed suppression for encoded, undecoded photons, and quantifies recovery under correct decoding $(\phi=0)$.
+Bandwidth matching
+For our rectangular spectrum of width $W$ in Eq. (9), the perfectly decoded case $(\phi=0)$ yields
+$\mid\mathcal{M}[0,\Delta]\mid^{2}\approx\frac{1}{N}\int_{-W/2}^{W/2}\frac{\gamma\operatorname{d}\omega}{\gamma^{2}/4+(\omega-\Delta)^{2}}$ (24)
+with $N=\int\operatorname{d}\omega\mid\xi(\omega)\mid^{2}$.
+This increases with $W/\gamma$ up to the point where the Lorentzian tails limit further gains, for a given $\gamma$ there is an optimal $W$ balancing spectral weight near resonance against total pulse energy.
+This explains why increasing code length $N_{0}$ (fixed $W$) does not improve excitation unless decoding removes the imposed phase structure.
+Our analysis uses a two-level, Markovian model with a single decay rate $\gamma$ and neglects recoil and multi-level structure.
+In free space, the achievable excitation also depends on the mode-matching parameter $\beta\leq 1$, which multiplies the overlap in Eq.
+(12).
+Non-Markovian reservoirs, structured continua, or multi-level $\Lambda$-systems can modify the optimal mode and the decoding strategy.
+Nonetheless, the overlap-based interpretation carries over: unitary spectral mode conversion preserves inner products, so matched decoding remains the key to recovering strong atom-photon coupling with spectrally encoded single photons.
+It is important to note that both encoding and decoding are unitary linear transformations in the single-photon Hilbert space, so they do not constitute a measurement and do not extract information about the encoded state.
+The process does not involve copying or amplifying a quantum state, rather, it merely rotates the state within its Hilbert space.
+Therefore, the no-cloning theorem, which forbids the creation of identical, independent copies of an arbitrary unknown state-remains fully respected.
+IV.2 Mismatch tolerance and phase-noise budget
+We assume small residual phase errors with zero mean and variance $\sigma_{\phi}^{2}$ over the occupied band. A second-order cumulant expansion gives
+$\mathcal{M}[\phi,\Delta]\approx\mathcal{M}[0,\Delta]e^{-\sigma_{\phi}^{2}/2}\implies P_{e}\approx P_{e}^{(0)}e^{-\sigma_{\phi}^{2}}$ (25)
+So a $10\%$ loss in $P_{e}$ corresponds to $\sigma_{\phi}\approx\sqrt{0.105}\approx 0.32$ rad rms across the in-band frequencies.
+This sets a concrete specification for pixel-to-pixel SLM (spatial light modulator) errors and spectral-phase drift in a fielded link.
+For a binary code with chips of width $\Omega=W/N_{0}$ and random sign errors of probability $p$, one obtains (Appendix B)
+$\langle e^{i\phi}\rangle=(1-p)(+1)+p(-1)=1-2p\implies P_{e}\approx P_{e}^{(0)}(1-2p)^{2}$ (26)
+so $p\leq 5\%$ keeps $P_{e}\geq 0.81P_{e}^{(0)}$.
+IV.3 Multiuser CDMA and cross-talk at an atomic node
+Consider $K$ simultaneous users with mutually assigned binary phase codes $\{\theta_{k}(\omega)\}$.
+The node intends to receive user $j$ and applies a decoding mask $D_{j}(\omega)=e^{i\theta_{j}(\omega)}$.
+The effective field is a sum $\sum_{k}\xi(\omega)e^{i[\theta_{j}(\omega)-\theta_{k}(\omega)]}$.
+The desired term has $\phi=0$; the $k\neq j$ terms contribute with code-difference masks.
+For random $\pm 1$ chip phases with length $N_{0}$, the expected value over random code assignments can be written as
+$\mathbb{E}\Big[\lvert\int\operatorname{d}\omega\xi(\omega)e^{i[\theta_{j}-\theta_{k}]}\chi_{A}(\omega)\rvert^{2}\Big]\propto\frac{1}{N_{0}}$ (27)
+i.e., cross-talk scales as $N_{0}^{-1}$.
+Hence, the single-node Signal-to-Interference Ratio satisfies $\text{SIR}\sim(K-1)^{-1}N_{0}$ (up to constants from spectral shaping and $\chi_{A}$).
+For practical guidance, we can keep $K\lesssim cN_{0}$ with $c<1$ (e.g., $c\sim 0.2-0.3)$ to maintain $\text{SIR}\gtrsim 3-5$ at the atomic interface.
+This formalizes the multiplexing benefit and the trade-off with code length observed in Section III.
+For more mathematical details, please refer to Appendix C.
+IV.4 Security and addressability via mode selectivity
+In the context of quantum communication, mode selectivity provides an inherent mechanism for both secure access and user addressability.
+The atomic system functions as a natural mode filter whose excitation probability $P_{e}$ depends quadratically on the spectral-temporal overlap between the incoming photon and the atomic response.
+When a photon arrives with a mismatched spectral-phase pattern, the destructive interference between frequency components suppresses the field amplitude at the atomic resonance.
+As demonstrated in Section III, the excitation probability then decreases rapidly with increasing code length $N_{0}$, yielding suppression factors that can reach several orders of magnitude for long encoded sequences.
+From a network standpoint, this behavior acts as a built-in physical layer of security: only photons whose spectral mode precisely matches the decoding mask $D(\omega)$ and the atomic resonance are capable of efficient excitation, while any unauthorized or misaligned signals are effectively rejected by the atom itself.
+This selectivity simultaneously enables addressability-each atomic node can be assigned a distinct decoding mask, allowing multiple users to share the same optical channel without classical routing.
+Consequently, the atom-photon interface serves as a passive, noise-free discriminator of spectral codes, enhancing both the privacy and scalability of quantum networks.
+IV.5 Co-design guidelines
+The quantitative analysis above suggests several practical guidelines for the co-design of spectral encoders, decoders, and atom-photon interfaces.
+First, bandwidth matching between the encoded photon spectrum and the atomic linewidth is essential.
+The excitation probability $P_{e}$ increases with the ratio $W/\gamma$ up to an optimum value where spectral weight begins to fall outside the Lorentzian response of the atom.
+We proved that for most single-transition systems, maximum excitation occurs for $W/\gamma\approx 1.39$ (Appendix D).
+Second, the code length $N_{0}$ should be chosen to balance multiplexing capacity and robustness.
+Longer codes improve the signal-to-interference ratio approximately as $N_{0}$ but they also increase sensitivity to phase errors.
+The tolerance analysis in Section IV B, shows that maintaining an rms phase error $\sigma_{\phi}\lesssim 0.3$ rad limits excitation loss to below $10\%$, implying a strict calibration requirement on the pulse shaper or spatial light modulator (SLM).
+Third, maximizing the coupling efficiency $\beta$ between the optical field and the atomic dipole remains the single most effective way to improve performance, as $P_{e}^{\max}\propto\beta\mid\mathcal{M}\mid^{2}$.
+High-NA (Numerical Aperture) free-space coupling, fiber-cavity, or nanophotonic waveguide geometries can substantially enhance $\beta$.
+Finally, spectral detuning $\Delta$ should be minimized or pre-compensated in the decoding phase $D(\omega)$ by introducing an appropriate linear phase term to center the photon spectrum on the atomic transition.
+Together, these guidelines translate the microscopic excitation model into an experimentally actionable design toolkit for encoded quantum-network links.
+IV.6 Experimental pathway
+The theoretical framework developed in this work can be validated on a variety of experimental platforms using existing photonic and atomic technologies. A representative configuration is as follows.
+A single-photon source, for example, spontaneous parametric down-conversion or quantum-dot emission spectrally shaped by a $4f$ pulse-shaping setup equipped with a programmable SLM that imposes the encoding phase pattern $\theta(\omega)$.
+The encoded photon propagates through an optical fiber channel to a receiving station implementing a decoding mask $D(\omega)=e^{i\tilde{\theta}(\omega)}$.
+The decoded photon then interacts with a two-level emitter such as a trapped atom, ion, or quantum dot coupled to a single-mode waveguide or high-NA free-space interface.
+The excitation probability is measured via fluorescence detection or by monitoring changes in transmission.
+Systematic variations of the code length $N_{0}$, phase-error amplitude $\sigma_{\phi}$, and decoding alignment allow direct experimental verification of (i) the exponential sensitivity $P_{e}\propto e^{-\sigma^{2}_{\phi}}$ predicted in Section IV B, (ii) the $N_{0}^{-1}$ scaling of cross-talk between multiple users (Section IV C), and (iii) the overall dependence of $P_{e}$ on $W/\gamma$ and $\beta$.
+Because all components-pulse shapers, fiber channels, and single emitters-are compatible with current laboratory technology, this platform-agnostic implementation offers a clear route for testing the theoretical predictions and for deploying spectral-encoded quantum communication protocols based on the principles developed in this study.
+V Conclusion
+We analyzed the interaction between a two-level atom and a spectrally encoded single-photon wave packet using the Heisenberg-Langevin approach.
+The results show that spectral phase encoding, while unitary, reshapes the photon’s temporal profile and reduces its peak intensity at the atom, thereby lowering the excitation probability compared with an unencoded pulse.
+The efficiency is governed by the temporal-mode overlap between the incident photon and the atom’s optimal absorption mode, incomplete decoding or phase errors directly reduce this overlap.
+From a network perspective, the atom behaves as a mode-selective and phase-sensitive receiver.
+Perfect conjugate decoding restores the original mode and strong coupling, whereas mismatched codes are naturally rejected, providing a built-in spectral filter for secure or multiplexed communication.
+Efficient operation requires careful control of spectral bandwidth, phase stability, mode matching, and detuning compensation.
+These parameters jointly determine the achievable overlap and thus the coupling efficiency.
+The framework presented here offers a compact description linking atomic excitation to a spectral-temporal overlap functional, supplying both a diagnostic for performance degradation under encoding and practical guidelines for implementing spectrally encoded photon-atom interfaces in scalable quantum networks.
 Acknowledgment
@@
 The data that support the findings of this study are available on request from the corresponding author
+Appendix A Derivation of the overlap bound in Eq. (12)
+We sketch a compact derivation of Eq. (12), which sets an upper bound on the atomic excitation probability for an arbitrary single-photon temporal mode $\xi(t)$.
+We consider a two-level atom coupled to a single input channel with coupling rate $\gamma$ and total decay rate $\Gamma_{\mathrm{tot}}$.
+Within the single-excitation manifold, the atomic excited-state amplitude $c_{e}(t)$ satisfies
+$\dot{c}_{e}(t)=-\frac{\Gamma_{\mathrm{tot}}}{2}c_{e}(t)+\sqrt{\gamma}\,\xi(t)$ (28)
+where we assume the atom is in the ground state in the remote past $c_{e}(-\infty)=0$.
+Eq. (A1) formal solution is
+$c_{e}(t)=\sqrt{\gamma}\!\int_{-\infty}^{t}e^{-(\Gamma_{\mathrm{tot}}/2)(t-\tau)}\,\xi(\tau)\,d\tau$ (29)
+At any time $t$, the instantaneous excitation probability is $P_{e}(t)=|c_{e}(t)|^{2}$.
+To obtain the largest possible excitation, we maximize $|c_{e}(t)|$ with respect to $\xi$ under the normalization
+$\int_{-\infty}^{+\infty}|\xi(t)|^{2}\,dt=1$ (30)
+Using the Cauchy-Schwarz inequality on Eq. (29) gives
+$|c_{e}(t)|^{2}$ $\leq\gamma\left(\int_{-\infty}^{t}e^{-\Gamma_{\mathrm{tot}}(t-\tau)}d\tau\right)\left(\int_{-\infty}^{t}|\xi(\tau)|^{2}d\tau\right)$
+$\leq\frac{\gamma}{\Gamma_{\mathrm{tot}}}\ \Big|\int_{-\infty}^{t}\sqrt{\Gamma_{\mathrm{tot}}}\,e^{-(\Gamma_{\mathrm{tot}}/2)(t-\tau)}\xi(\tau)\,d\tau\Big|^{2}$ (31)
+The inequality is saturated when $\xi(\tau)$ is proportional to the kernel of the convolution, i.e.
+$\xi_{\mathrm{opt}}(t)=\sqrt{\Gamma_{\mathrm{tot}}}\,e^{+(\Gamma_{\mathrm{tot}}/2)t}\Theta(-t)$ (32)
+which is the time-reversed spontaneous-emission mode. If we denote $\gamma=\Gamma_{\mathrm{tot}}$, we obtain Eq. (11) which is exactly the optimum (normalaized) temporal profile.
+Defining the coupling efficiency $\beta=\gamma/\Gamma_{\mathrm{tot}}$ and normalizing $\xi_{\mathrm{opt}}$ so that $\int|\xi_{\mathrm{opt}}(t)|^{2}dt=1$, we obtain
+$P_{e}^{\max}\;=\;\beta\,\big|\!\langle\xi\mid\xi_{\mathrm{opt}}\rangle\!\big|^{2},\qquad\langle\xi\mid\xi_{\mathrm{opt}}\rangle=\int_{-\infty}^{+\infty}\xi^{*}(t)\xi_{\mathrm{opt}}(t)\,dt$ (33)
+Equation (33) is Eq.(12) in the main text and expresses the maximal excitation probability as the product of the spatial-mode coupling $\beta$ and the temporal-mode overlap between the incident photon and the optimal rising exponential mode.
+Appendix B Effect of binary chip-flip errors on excitation
+We partition the occupied band into $N_{0}$ chips of width $\Omega=W/N_{0}$.
+After ideal decoding, the desired field adds coherently across chips with complex weights $\{a_{n}\}_{n=1}^{N_{0}}$, so the (error-free) excitation amplitude is $A^{(0)}=\sum_{n}a_{n}$ and $P_{e}^{(0)}\propto|A^{(0)}|^{2}$.
+Now suppose a binary phase error model: on each chip $n$, a random sign flip occurs with probability $p$, i.e.
+$\phi_{n}=\begin{cases}0&\text{w.p. }1-p,\\ \pi&\text{w.p. }p,\end{cases}\qquad X_{n}\equiv e^{i\phi_{n}}\in\{+1,-1\}$ (34)
+independently across $n$. Then
+$\langle e^{i\phi_{n}}\rangle=\langle X_{n}\rangle=(1-p)(+1)+p(-1)=1-2p$ (35)
+With errors, the excitation amplitude is
+$A=\sum_{n=1}^{N_{0}}a_{n}X_{n},\qquad\Rightarrow\qquad\langle A\rangle=(1-2p)\sum_{n=1}^{N_{0}}a_{n}=(1-2p)\,A^{(0)}$ (36)
+Exact interference-power expression.
+Because $X_{n}$ are independent, $X_{n}^{2}\equiv 1$, and $\langle X_{n}X_{m}\rangle=\langle X_{n}\rangle\langle X_{m}\rangle=(1-2p)^{2}$ for $n\neq m$,
+$\left\langle|A|^{2}\right\rangle$ $=\sum_{n=1}^{N_{0}}|a_{n}|^{2}\langle X_{n}^{2}\rangle+\sum_{\begin{subarray}{c}n,m=1\\ n\neq m\end{subarray}}^{N_{0}}a_{n}a_{m}^{*}\langle X_{n}X_{m}\rangle$
+$=\sum_{n=1}^{N_{0}}|a_{n}|^{2}+(1-2p)^{2}\sum_{\begin{subarray}{c}n,m=1\\ n\neq m\end{subarray}}^{N_{0}}a_{n}a_{m}^{*}$
+$=(1-2p)^{2}\Big|\sum_{n=1}^{N_{0}}a_{n}\Big|^{2}\;+\;\big[1-(1-2p)^{2}\big]\sum_{n=1}^{N_{0}}|a_{n}|^{2}$
+$=(1-2p)^{2}|A^{(0)}|^{2}\;+\;4p(1-p)\sum_{n=1}^{N_{0}}|a_{n}|^{2}$ (37)
+Thus the exact mean excitation probability is
+$\langle P_{e}\rangle\ \propto\ (1-2p)^{2}\,|A^{(0)}|^{2}\;+\;4p(1-p)\sum_{n=1}^{N_{0}}|a_{n}|^{2}$ (38)
+Coherent-sum (large-$N_{0}$) approximation.
+When the decoded chips add coherently (e.g., $a_{n}\approx a$) we have $|A^{(0)}|^{2}=N_{0}^{2}|a|^{2}$ and $\sum_{n}|a_{n}|^{2}=N_{0}|a|^{2}$, so the second term in (37) is smaller by a factor $\sim 1/N_{0}$:
+$\frac{4p(1-p)\sum_{n}|a_{n}|^{2}}{(1-2p)^{2}|A^{(0)}|^{2}}=\frac{4p(1-p)}{(1-2p)^{2}}\cdot\frac{1}{N_{0}}\ \ll\ 1\qquad(N_{0}\ \text{large})$
+Hence
+$P_{e}\ \approx\ P_{e}^{(0)}\,(1-2p)^{2}$ (39)
+which is the stated result in Eq. (26). Eq. (39) follows directly from the mean chip factor (35) and the dominance of the coherent term for large $N_{0}$; the exact correction is given by (37).
+Appendix C Expectations over random code assignments and SIR
+We split the occupied band into $N_{0}$ chips $\{\omega_{n}\}_{n=1}^{N_{0}}$.
+Each user $k$ employs a phase code $\theta_{k}(\omega_{n})$.
+We average over i.i.d.
+random codes drawn from the binary ensemble
+$\mathcal{E}_{\mathrm{bin}}:\quad\theta_{k}(\omega_{n})\in\{0,\pi\}\ \text{i.i.d. with}\ \Pr[\theta_{k}(\omega_{n})=0]=\Pr[\theta_{k}(\omega_{n})=\pi]=\tfrac{1}{2}$ (40)
+and denote expectations by $\mathbb{E}_{\mathrm{codes}}[\cdot]$. For two users $j\neq k$, we define the chipwise code-difference factor
+$X_{n}\equiv e^{i[\theta_{j}(\omega_{n})-\theta_{k}(\omega_{n})]}\in\{+1,-1\}$ (41)
+Under $\mathcal{E}_{\mathrm{bin}}$, $\{X_{n}\}$ are i.i.d. Rademacher variables with
+$\mathbb{E}[X_{n}]=0,\qquad\mathbb{E}[X_{n}^{2}]=1$ (42)
+The normalized discrete code correlation that appears in overlap integrals is
+$C_{jk}\ \equiv\ \frac{1}{N_{0}}\sum_{n=1}^{N_{0}}X_{n}\ =\ \frac{1}{N_{0}}\sum_{n=1}^{N_{0}}e^{i[\theta_{j}(\omega_{n})-\theta_{k}(\omega_{n})]}$ (43)
+By linearity and independence,
+$\mathbb{E}_{\mathrm{codes}}[C_{jk}]$ $=\frac{1}{N_{0}}\sum_{n=1}^{N_{0}}\mathbb{E}[X_{n}]=0$ (44)
+$\mathrm{Var}_{\mathrm{codes}}(C_{jk})$ $=\mathbb{E}\!\left[\left|\frac{1}{N_{0}}\sum_{n=1}^{N_{0}}X_{n}\right|^{2}\right]=\frac{1}{N_{0}^{2}}\sum_{n=1}^{N_{0}}\mathbb{E}[X_{n}^{2}]=\frac{1}{N_{0}}$ (45)
+Hence $C_{jk}\xrightarrow{\mathbb{P}}0$ as $N_{0}\to\infty$, and typical cross-talk scales as $N_{0}^{-1/2}$ in amplitude (or $N_{0}^{-1}$ in power).
+Let $a_{n}$ denote the deterministic complex amplitude on chip $n$ for the desired user $j$ after decoding (it can collect spectral weight, atomic response, and any fixed factors).
+The interference at the node from another user $k\neq j$ is
+$I_{jk}\ \equiv\ \sum_{n=1}^{N_{0}}a_{n}\,X_{n}$ (46)
+Then
+$\mathbb{E}_{\mathrm{codes}}[I_{jk}]=0,\qquad\mathbb{E}_{\mathrm{codes}}\!\big[\,|I_{jk}|^{2}\,\big]=\sum_{n=1}^{N_{0}}|a_{n}|^{2}$ (47)
+since cross terms vanish by independence and $\mathbb{E}[X_{n}X_{m}]=0$ for $n\neq m$. With K1K-1 independent interfering users,
+$\mathbb{E}_{\mathrm{codes}}\!\left[\left|\sum_{k\neq j}I_{jk}\right|^{2}\right]=(K-1)\sum_{n=1}^{N_{0}}|a_{n}|^{2}$ (48)
+The desired (decoded) field amplitude for user $j$ is
+$S_{j}\ \equiv\ \sum_{n=1}^{N_{0}}a_{n}$ (49)
+so a natural signal-to-interference ratio (SIR) is
+$\mathrm{SIR}\ \equiv\ \frac{|S_{j}|^{2}}{\mathbb{E}_{\mathrm{codes}}\!\left[\sum_{k\neq j}|I_{jk}|^{2}\right]}=\frac{\left|\sum_{n=1}^{N_{0}}a_{n}\right|^{2}}{(K-1)\sum_{n=1}^{N_{0}}|a_{n}|^{2}}$ (50)
+For flat chips $a_{n}=a$,
+$|S_{j}|^{2}=|a|^{2}N_{0}^{2},\qquad\sum_{n}|a_{n}|^{2}=|a|^{2}N_{0}\quad\Rightarrow\quad\ \mathrm{SIR}=\dfrac{N_{0}}{K-1}\$(51)
+Appendix D Optimal spectral bandwidth for maximum excitation
+The excitation probability of a two-level atom driven by a single photon with normalized spectral amplitude $\xi(\omega)$ is proportional to the squared overlap between $\xi(\omega)$ and the atomic susceptibility $\chi_{A}(\omega)$,
+$P_{e}(W)\ \propto\ \Big|\int_{-\infty}^{\infty}\xi(\omega)\,\chi_{A}(\omega)\,d\omega\Big|^{2}$ (52)
+Considering a rectangular photon spectrum of total width $W$ centered at the atomic transition frequency $\omega_{A}$, and the frequency-domain response $\chi_{A}(\omega)$ in Eq. (20), Eq. (52) gives
+$\mathcal{M}(W)=\frac{\sqrt{\gamma}}{\sqrt{W}}\int_{-W/2}^{W/2}\frac{d\omega}{\gamma/2-i\omega}=\frac{2\sqrt{\gamma}}{\sqrt{W}}\tan^{-1}\!\left(\frac{W}{\gamma}\right)$ (53)
+where $\mathcal{M}(W)$ represents the (dimensionless) spectral overlap amplitude. The excitation probability then scales as
+$P_{e}(W)\ \propto\ \frac{\gamma}{W}\left[\tan^{-1}\!\left(\frac{W}{\gamma}\right)\right]^{2}$ (54)
+The behavior of Eq. (54) can be analyzed in the two limiting regimes:
+$P_{e}(W)$ $\propto W,\qquad$ $W\ll\gamma$ (55)
+$P_{e}(W)$ $\propto\frac{1}{W},\qquad$ $W\gg\gamma$ (56)
+Therefore, $P_{e}(W)$ must reach a maximum for an intermediate bandwidth. Setting $\partial P_{e}/\partial W=0$ in Eq. (54) yields a transcendental equation
+$\tan^{-1}\!\left(W/\gamma\right)=2(W/\gamma)/(1+(W/\gamma)^{2})$ (57)
+whose numerical solution is $W/\gamma\simeq 1.39$.
+Physically, this condition corresponds to matching the photon’s spectral width to the atomic linewidth: if $W\ll\gamma$, the pulse is too long in time and couples weakly, whereas for $W\gg\gamma$ much of the photon energy lies off-resonance and is not absorbed.
+The optimal intermediate regime ensures efficient temporal and spectral overlap between the encoded photon and the atomic transition.